Find the marginal average cost function if cost and revenue are given by C(x) = 158 + 3.6x and R(x) = 4x – 0.03x2. The marginFind the marginal average cost function if cost and revenue are given by ​C(x)=158+3.6x and R(x)=4x−0.03x2. The marginal average cost function is C′(x)= nothing.
Answer #1

C(x)=158+3.6x
Average cost function = 158/x + 3.x/x = 158x^-1 + 3.6
Taking derivating we get:

Marginal average cost function = -158x^-2 or -158/x^2